OpenStudy (anonymous):

Evaluate the following expression. 2(x + 9) for x = 4

OpenStudy (anonymous):

2(4+9)=26

OpenStudy (anonymous):

how did you get 26?

OpenStudy (anonymous):

just insert 4 for the x, then you get 2(4+9) = 2(13) = 2*13 = 26